What is the value of the equilibrium constant at 25 °C for a reaction that has a standard free-energy change, ΔG°, equal to 180 kJ?
Answer #1

T= 25.0 oC

= (25.0+273) K

= 298 K

ΔGo = 180 KJ/mol

ΔGo = 180000 J/mol

use:

ΔGo = -R*T*ln Kc

180000 = - 8.314*298.0* ln(Kc)

ln Kc = -72.6518

Kc = 2.804*10^-32

Answer: 2.80*10^-32